Am given Qd=40-49 % As=-16+4p
Solve algebraical and determine the Equilibrium price and Quantity.
1
Expert's answer
2020-09-29T09:10:53-0400

In equilibrium Qd = Qs, so:

"40 - 49p = -16 + 4p,"

53p = 56,

p = 1.057,

"Q = 40 - 49\u00d71.057 = -11.77."